# On the Reachability of 3–Dimensional Paths with a Prescribed Curvature Bound

### v2

@@ 2.3 Reachability of Dubins car
Apart from the Markov–Dubins problem, there had been several attempts to obtain the reachability set of planar curves with prescribed curvature bound. To facilitate subsequent discussions, we introduce the following notations. For each time $t$, we will denote the reachability set of the dynamical system Eq. (6) as $\mathcal{G}_{2}(t)$. By defining a projection map $\pi_{2}:(x^{1},x^{2},\theta)\mapsto(x^{1},x^{2})$, the position reachability set is indicated as $\pi_{2}\left(\mathcal{G}_{2}(t)\right)$. Similarly, we denote the reachability set of the dynamical system Eq. (9) as $\mathcal{G}_{3}(t)$. By defining a projection map $\pi_{3}:(\boldsymbol{x},\boldsymbol{y})\mapsto\boldsymbol{x}$, the position reachability set is indicated as $\pi_{3}\left(\mathcal{G}_{3}(t)\right)$.
@@
Unlike Cockayne and Hall (1975) which relied on geometric arguments, the reachability analysis of the Dubins car with consideration of directional components, or the identification of $\mathcal{G}_{2}(t)$, in Patsko et al. (2003) leveraged Theorem 3. It was proved that every boundary point of $\mathcal{G}_{2}(t)$ can be reached by trajectories of CSC, CCC, or their subsegments. Although Theorem 3 in principle does not provide complete determination of the boundary, within the specific dynamics of the Dubins car in Eq. (6) and small–time $\leq\frac{3.63\pi}{\kappa_{M}}$, Theorem 3 was sufficient for construction of the boundary of $\mathcal{G}_{2}(t)$. This was because for small times, the surfaces generated by collecting the endpoints of the trajectories of CSC, CCC, and their subsegments altogether generated a single surface enclosing a simply–connected set in $\mathbb{R}^{3}$. (Figs. 8 and 9 in Patsko et al. (2003)) The complete determination of $\mathcal{G}_{2}(t)$ for general $t$ was finally achieved in the follow–up study of Patsko and Fedotov (2022).
@@ 3.1 Connection with time optimal problems
In this subsection, we present the internal connection between reachability and time optimal problems through PMP. Let us consider the following OCP which final time is considered as a cost, denoted as $\textbf{P}_{\boldsymbol{x}_{f},\phi}$.
@@
where the Hamiltonian is defined as $\mathcal{H}(p_{0},\boldsymbol{p},\boldsymbol{x},\boldsymbol{u})\equiv\boldsymbol{p}^{T}f(\boldsymbol{x},\boldsymbol{u})+p_{0}\phi$. Here, a pair $\{p^{*}_{0},\boldsymbol{p}^{*}\}$ is said to be nontrivial if $p^{*}_{0}\neq 0$ or $\boldsymbol{p}^{*}$ is nonvanishing as a covector on $M$. From here on, any triplet $\{\boldsymbol{x}^{*},\boldsymbol{u}^{*},\boldsymbol{p}^{*}\}$ satisfying the PMP conditions of an OCP on an interval $[0,t_{f}^{*}]$ is referred to extremal. Subsequently, the set of time optimal problems and their extremals are defined as follows.
Definition 8.
+$\begin{split}&\mathbb{P}^{+}\equiv\{\textbf{P}_{\boldsymbol{x}_{f},\phi}:\boldsymbol{x}_{f}\in\mathbb{R}^{n_{\boldsymbol{x}}},\enspace\phi>0\}\\ &\mathbb{P}^{-}\equiv\{\textbf{P}_{\boldsymbol{x}_{f},\phi}:\boldsymbol{x}_{f}\in\mathbb{R}^{n_{\boldsymbol{x}}},\enspace\phi<0\}\\ &\mathbb{P}\equiv\mathbb{P}^{+}\cup\mathbb{P}^{-}\\ &\tilde{S}^{*}(t_{f})\equiv\{\{\boldsymbol{x}^{*},\boldsymbol{u}^{*},\boldsymbol{p}^{*}\}:\text{extremals of }\mathbb{P}\text{ s.t. }t_{f}^{*}=t_{f}\}\end{split}$
[…]
Theorem 9.
$S^{*}(t_{f})=\tilde{S}^{*}(t_{f})$.
{pf*}
Proof. For some solution $\{\boldsymbol{x}^{*},\boldsymbol{u}^{*},\boldsymbol{p}^{*}\}\in S^{*}(t_{f})$, define $\tilde{\boldsymbol{x}}^{*}=\boldsymbol{x}^{*}$, $\tilde{\boldsymbol{u}}^{*}=\boldsymbol{u}^{*}$, and $\tilde{\boldsymbol{p}}^{*}=\boldsymbol{p}^{*}$ on $[0,t_{f}]$. Then the costate differential and pointwise maximum conditions of $\textbf{P}_{\boldsymbol{x}_{f},\phi}$ are readily satisfied by such choice of $\{\tilde{\boldsymbol{x}}^{*},\tilde{\boldsymbol{u}}^{*},\tilde{\boldsymbol{p}}^{*}\}$. Nontriviality condition follows from nontriviality of $\boldsymbol{p}^{*}$ by itself. Since the system is autonomous, the optimal Hamiltonian $\langle\boldsymbol{p}^{*},f\rangle=\langle\tilde{\boldsymbol{p}}^{*},f\rangle$ is constant over time. Then existence of $\tilde{p}_{0}^{*}\geq 0$ and $\phi\neq 0$ such that $\tilde{p}_{0}^{*}\phi=-\langle\tilde{\boldsymbol{p}}^{*},f\rangle\in\mathbb{R}$ is obvious. Therefore, such $\{\tilde{\boldsymbol{x}}^{*},\tilde{\boldsymbol{u}}^{*},\tilde{\boldsymbol{p}}^{*},\tilde{p}_{0}^{*}\}$ solves the PMP condition of an OCP in $\mathbb{P}$ and hence, $\{\boldsymbol{x}^{*},\boldsymbol{u}^{*},\boldsymbol{p}^{*}\}=\{\tilde{\boldsymbol{x}}^{*},\tilde{\boldsymbol{u}}^{*},\tilde{\boldsymbol{p}}^{*}\}\in\tilde{S}^{*}(t_{f})$. Thus, $S^{*}(t_{f})\subseteq\tilde{S}^{*}(t_{f})$.
Conversely, suppose $\{\tilde{\boldsymbol{x}}^{*},\tilde{\boldsymbol{u}}^{*},\tilde{\boldsymbol{p}}^{*}\}\in\tilde{S}^{*}(t_{f})$, where $\{\tilde{p}_{0}^{*},\tilde{\boldsymbol{p}}^{*}\}$ is the corresponding adjoint response. Then the costate differential and pointwise maximum conditions in Eqs. (3) and (4) are met by choosing $\{\boldsymbol{x}^{*},\boldsymbol{u}^{*},\boldsymbol{p}^{*}\}=\{\tilde{\boldsymbol{x}}^{*},\tilde{\boldsymbol{u}}^{*},\tilde{\boldsymbol{p}}^{*}\}$. If $\tilde{\boldsymbol{p}}^{*}$ is trivial (i.e., $\tilde{\boldsymbol{p}}^{*}$ vanishes as covector on $M$), then $\langle\tilde{\boldsymbol{p}}^{*},f(\tilde{\boldsymbol{x}}^{*},\tilde{\boldsymbol{u}}^{*})\rangle=0$ since $f(\tilde{\boldsymbol{x}}^{*},\tilde{\boldsymbol{u}}^{*})$ is a tangent vector of $M$ at $\tilde{\boldsymbol{x}}^{*}$. Then everywhere vanishing Hamiltonian condition implies $\tilde{p}_{0}^{*}=0$, which contradicts the nontriviality condition. Thus, $\tilde{\boldsymbol{p}}^{*}$ is nontrivial by itself and nontriviality of $\boldsymbol{p}^{*}$ follows. This proves that $\{\tilde{\boldsymbol{x}}^{*},\tilde{\boldsymbol{u}}^{*},\tilde{\boldsymbol{p}}^{*}\}=\{\boldsymbol{x}^{*},\boldsymbol{u}^{*},\boldsymbol{p}^{*}\}\in S^{*}(t_{f})$. ∎ Theorem 9 generalizes the one–sided inclusion relationship identified in Lewis (2006) that time optimal trajectories lie on the boundary of the reachability set. The inclusion in the converse direction, which asserts that any trajectory reaching the boundary of the reachability set corresponds to an extremal of a time optimal problem, is made possible by considering the extremals instead of optimal solutions of OCPs. Subsequently, this enables to utilize the existing results on the time optimal problems to obtain useful conclusions on reachability sets. The following remark is the core idea of the reachability analysis of 3D curves with prescribed curvature bound in the subsequent discussions.
Remark 10.
Once the extremals of $\mathbb{P}^{+}$ (resp. $\mathbb{P}^{-}$) are known on $[0,t_{f}]$, Theorem 9 implies that $E^{*}(t_{f})$ can be determined by solving the PMP conditions of $\mathbb{P}^{-}$ (resp. $\mathbb{P}^{+}$). During this procedure, given that the only difference between the PMP conditions of $\mathbb{P}^{+}$ and $\mathbb{P}^{-}$ is the sign of $\phi$, any conclusion that is independent of the sign of $\phi$ can be utilized when solving the PMP conditions for the counterpart.
@@ 3.2 Reachability of 3D curves with prescribed curvature bound
In this subsection, we present the reachability analysis of 3D curves with prescribed curvature bound. The analysis is organized into two parts: first, we provide an implicit description of $\mathcal{G}_{3}(t_{f})$ by proving that every boundary point can be reached by curves belonging to the classes CCC, CSC, their subsegments, or H. Second, we present the complete description of $\pi_{3}\left(\mathcal{G}_{3}(t_{f})\right)$.
@@ 3.2.1 Implicit description of the reachability set
Proof. Proof of this theorem is largely an application of Theorem 3 to the system Eq. (9). For sake of completeness, we mention that Eq. (9) satisfies Assumption 1 because it is a smooth control affine system with compact and convex restraint set, $\mathbb{B}^{3}$. (Recall that the maps $\boldsymbol{w}\mapsto\boldsymbol{y}\times\boldsymbol{w}$ correspond to skew–symmetric linear transformations of $\mathbb{R}^{3}$.) Theorem 9 further implies that identifying $E^{*}(t_{f})$ is equivalent with identifying the endpoints of the extremals of problems in $\mathbb{P}$.
As introduced previously, Sussmann (1995) provides a detailed description of the extremals of the minimum time problems described by Eqs. (8), (9), and (10). In other words, the extremals of $\mathbb{P}^{-}$ problems are studied. In this regard, some portions of the proof of this theorem inherently have considerable overlap with the proof of Theorem 1 in Sussmann (1995). As outlined in Remark 10, this is the virtue of Theorem 9: once the extremals of the minimum time problems are analyzed, it suffices to redo the proof with a gentle modification to find the extremals of the maximum time problems (i.e., $\mathbb{P}^{+}$ problems) to find $E^{*}(t_{f})$. Therefore, during the proof of this theorem and the subsequent ones, the notations we define will follow the same ones used in the proof of Theorem 1 in Sussmann (1995). The modifications required are threefold and can be summarized as follows:
  • we ought to extend the analysis to consider $\mathbb{P}^{+}$ problems as well,
  • the cycle argument should not be used,
  • we can exclude some points in $E^{*}(t_{f})$ such that there exists an alternative trajectory reaching the point that only allows trivial adjoint response.
The proof is organized as follows. First, in the ‘Basic setup’ part, we will introduce the basic notations and summarize the developments from Sussmann (1995) that will be used in the subsequent part. Since the three modifications previously mentioned are irrelevant with this step, this part is essentially contained in Sussmann (1995). In the subsequent Part 2, we prove items (1) and (2) of this theorem with the three modifications. In the last Part 3, we prove the remaining items (3), (4), and (5).
@@
In terms of the notations used in the previous sections, $(\boldsymbol{\lambda},\boldsymbol{\mu})$ corresponds to the adjoint variable $-\boldsymbol{p}$ and $\nu$ corresponds to $-p_{0}\phi$. From here on, we will denote $\boldsymbol{\Xi}=(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\lambda},\boldsymbol{\mu},\nu,\boldsymbol{w})$ satisfying Eqs. (9) and (15) by lifted controlled trajectories, and lifted controlled arcs if it is defined on a compact interval. A lifted controlled trajectory is said to be $\mathcal{H}$–minimizing if $\mathcal{H}\left(\boldsymbol{\Xi}(t)\right)=\min\{\mathcal{H}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\lambda},\boldsymbol{\mu},\nu,\boldsymbol{v}):\boldsymbol{v}\in\mathbb{B}^{3}\}$ for a.e. $t$.
@@
    $\boldsymbol{\Xi}$ is $\mathcal{H}$–minimizing,
@@
These are merely a restatement of the PMP conditions of $\textbf{P}_{\boldsymbol{x}_{f},\phi}$ with arbitrary $\boldsymbol{x}_{f}$ and $\phi\neq 0$. Indeed, when studying the minimum time problem, or $\mathbb{P}^{-}$ problems, in Sussmann (1995), $\nu\geq 0$ condition was present. This condition is omitted here because we consider $\mathbb{P}^{+}$ problems as well.
@@
Dependency of $\boldsymbol{\lambda}$ on $t$ is ignored since Eq. (15) implies that $\boldsymbol{\lambda}$ is constant in $t$. In Sussmann (1995), this formulation of the nontriviality condition—rather than the “naïve” version that requires $\left(\boldsymbol{\lambda},\boldsymbol{\mu}\right)\neq 0$—played a crucial role. This is because the maximum principle with the latter nontriviality condition provided no useful information, as every trajectory was extremal. This point was also outlined in Remark 3.2 in Sussmann (1995). Under the nontriviality condition in Eq. (16), the remaining part of the proof of this theorem demonstrates how maximum principle provides nontrivial conditions to sort out extremals.
@@
Our proof follows the similar step, but with the three aforementioned modifications. For the rest of part 1, we introduce the major conclusions—primarily on the structure of nice trajectories—made in Sussmann (1995) that were solely grounded on the PMP conditions. Theorem 9 asserts that these conclusions can be applied analogously to our proof. The summary is as follows:
@@
If $C\neq 0$, Eq. (22) can be modified in terms of $\tau$, which becomes exactly Eq. (11) with $\zeta=\frac{2\nu}{\sqrt{|C|}}$. Since nice trajectories are globally defined in this case, there is no need of further analysis how these connect up with trajectories such that $\boldsymbol{W}=0$. Additional remark on $C\neq 0$ case is that the condition of $\nu\geq 0$ was present in Sussmann (1995) because $\mathbb{P}^{-}$ problems were considered. Hence, the ‘H’ curves were confined to the ones with $\zeta\geq 0$ but here we allow $\zeta<0$ case as well.
@@
In this part, we study how nice trajectories connect up with trajectories such that $\boldsymbol{W}=0$ when it is not globally defined. This is essentially a process of identifying $S^{*}(t_{f})$ and $E^{*}(t_{f})$, but we alongside exclude some trajectories that have an alternative trajectory reaching the same endpoint but only allows trivial adjoint response.
Let us commence by considering a lifted controlled arc $\boldsymbol{\Xi}$ defined on an interval $I=[0,t_{f}]$ satisfying the PMP conditions. We define $L=\{t\in I:\boldsymbol{W}(t)\neq 0\}$ and its complement in $I$, $Q=I\setminus L=\{t\in I:\boldsymbol{W}(t)=0\}$. Since $\|\boldsymbol{W}\|=\|\boldsymbol{y}\times\boldsymbol{\mu}\|$ is obviously continuous, its inverse image of the open set $\mathbb{R}\setminus\{0\}$, $L$, is open in $I$.
@@
First, if $\|\boldsymbol{\lambda}\|<|\nu|$, then Eq. (24) is a contradiction. Thus, $\boldsymbol{W}\neq 0$ and $\boldsymbol{\Xi}$ is a single piece of nice arc if it exists.
If $\|\boldsymbol{\lambda}\|>|\nu|$, then $\dot{\boldsymbol{W}}(t_{1})\neq 0$ for all $t_{1}\in Q$. So each $t_{1}$ is an isolated point of $Q$ and $Q$ is a discrete set. This implies that $\boldsymbol{\Xi}$ consists of concatenation of number of nice trajectories that are connected up with each other at points such that $\boldsymbol{W}=0$.
If $\|\boldsymbol{\lambda}\|=|\nu|$ then the zeros of $\boldsymbol{W}$ are not necessarily isolated and $\boldsymbol{\Xi}$ could consist of concatenations of nice trajectories with trajectories such that $\boldsymbol{W}\equiv 0$.
@@
Since $\boldsymbol{W}$ is differentiable, Taylor’s Theorem implies that for $t_{1}\in Q$, $\boldsymbol{W}(t)=\dot{\boldsymbol{W}}(t_{1})(t-t_{1})+o(|t-t_{1}|)$ and $\boldsymbol{w}(t)=sgn(t-t_{1})\frac{\dot{\boldsymbol{W}}(t_{1})}{\|\dot{\boldsymbol{W}}(t_{1})\|}+o(1)$ for $t\neq t_{1}$ near $t_{1}$. Therefore, the left and right side limits of $\boldsymbol{w}$ only change signs at $t_{1}$. This implies that the osculating plane of the two circular arcs being concatenated at $t_{1}$ coincides and their unit normal vectors are in the opposite directions. Hence, the state trajectory $\boldsymbol{x}(t)$ is a planar curve, concatenation of countable circular arcs.
Moreover, $\boldsymbol{W}=0$ at each point of concatenation. Thus, $\boldsymbol{\Xi}_{i}$ for $12\leq i\leq N-$ 1 correspond to maximal solutions of Eq. (22). Here, $\boldsymbol{\Xi}_{1}$ and $\boldsymbol{\Xi}_{N}$ are not necessarily maximal because $\boldsymbol{W}(0)$ and $\boldsymbol{W}(t_{f})$ are not necessarily 00.
@@
If $\nu=-\|\boldsymbol{\lambda}\|$, then it is easy to see that $\|\boldsymbol{W}\|=\|\boldsymbol{\lambda}\|\cos(t-t_{0})+\nu$ can be nonnegative only at isolated points. Hence, no lifted controlled trajectory $\boldsymbol{\Xi}$ on a nontrivial interval can exist if $\nu=-\|\boldsymbol{\lambda}\|$.
If $\nu=\|\boldsymbol{\lambda}\|$, then maximal solutions of Eq. (22) are defined on open intervals of length $2\pi$. Since $L$ is open in $I$, it can be expressed as $L=\bigcup\limits_{J\in\mathcal{J}}J$ where $\mathcal{J}$ is a countable set of disjoint open subintervals of $I$. Then for any $J\in\mathcal{J}$ that does not contain 00 and $t_{f}$, $\boldsymbol{W}=0$ at its boundary points. Thus, such $J$ has a length of $2\pi$ and corresponds to the domain of a maximal nice trajectory. These maximal nice trajectories correspond to a single cycle. If the domain contains 00 or $t_{f}$, then corresponding $\varphi$ does not have to be maximal, and naturally, such nice trajectory is a circular arc of length $\leq 2\pi$.

### v3

@@ 2.3 Reachability of Dubins car
Apart from the Markov-Dubins problem, there had been several attempts to obtain the reachability set of planar curves with prescribed curvature bound. To facilitate subsequent discussions, we introduce the following notations. For each time $t$, we will denote the reachability set of the dynamical system Eq. (6) as $\mathcal{G}_{2}(t)$. By defining a projection map $\pi_{2}:(x^{1},x^{2},\theta)\mapsto(x^{1},x^{2})$, the position reachability set is indicated as $\pi_{2}\left(\mathcal{G}_{2}(t)\right)$. Similarly, we denote the reachability set of the dynamical system Eq. (9) as $\mathcal{G}_{3}(t)$. By defining a projection map $\pi_{3}:(\boldsymbol{x},\boldsymbol{y})\mapsto\boldsymbol{x}$, the position reachability set is indicated as $\pi_{3}\left(\mathcal{G}_{3}(t)\right)$.
@@
Unlike Cockayne and Hall (1975) which relied on geometric arguments,  Patsko et al. (2003) leveraged Theorem 3 to analyze the reachability set with consideration of directional components, or $\mathcal{G}_{2}(t)$. It was proved that every boundary point of $\mathcal{G}_{2}(t)$ can be reached by trajectories of CSC, CCC, or their subsegments. Although Theorem 3 in principle does not provide complete determination of the boundary, within the specific dynamics of the Dubins car in Eq. (6) and small-time $\leq\frac{3.63\pi}{\kappa_{M}}$, Theorem 3 was sufficient for construction of the boundary of $\mathcal{G}_{2}(t)$. This was because for small times, the surfaces generated by collecting the endpoints of the trajectories of CSC, CCC, and their subsegments altogether generated a single surface enclosing a simply-connected set in $\mathbb{R}^{3}$. (Figs. 8 and 9 in Patsko et al. (2003)) The complete determination of $\mathcal{G}_{2}(t)$ for general $t$ was finally achieved in the follow-up study of Patsko and Fedotov (2022).
@@ 3.1 Connection with time optimal problems
In this subsection, we present how the results in Sussmann (1995) can be related with reachability analysis. Let us consider the following OCP which final time is considered as a cost, denoted as $\textbf{P}_{\boldsymbol{x}_{f},\phi}$.
@@
where Hamiltonian is defined as $\mathcal{H}(p_{0},\boldsymbol{p},\boldsymbol{x},\boldsymbol{u})\equiv\boldsymbol{p}^{T}f(\boldsymbol{x},\boldsymbol{u})+p_{0}\phi$. A pair $\{p^{*}_{0},\boldsymbol{p}^{*}\}$ is said to be nontrivial if $p^{*}_{0}\neq 0$ or $\boldsymbol{p}^{*}$ is nonvanishing as a covector on $M$. Any response satisfying the PMP conditions on an interval $[0,t_{f}^{*}]$ will be referred to as maximal response and the corresponding controller as maximal controller. The triplet $\{\boldsymbol{x}^{*},\boldsymbol{u}^{*},\boldsymbol{p}^{*}\}$ is referred to as maximal.
It is easy to notice that the costate differential and pointwise maximum conditions of $\textbf{P}_{\boldsymbol{x}_{f},\phi}$ problems are equivalent with Eqs. (3) and (4). Moreover, from everywhere vanishing Hamiltonian condition, vanishing $\boldsymbol{p}^{*}$ as a covector of $M$ implies $p_{0}^{*}=0$. Thus, $\boldsymbol{p}^{*}$ has to be nontrivial by itself. This implies that a maximal of a $\textbf{P}_{\boldsymbol{x}_{f},\phi}$ problem becomes a nontrivial solution to Eqs. (3) and (4). Conversely, $\boldsymbol{p}^{*}(t)^{T}f(\boldsymbol{x}^{*}(t),\boldsymbol{u}^{*}(t))$ remains constant over time since $f$ is an autonomous system, and existence of $p_{0}^{*}$ satisfying the everywhere vanishing Hamiltonian condition follows thereby. Consequently, collecting all nontrivial solutions of Eqs. (3) and (4) is equivalent with finding all maximals of $\textbf{P}_{\boldsymbol{x}_{f},\phi}$ problems. In other words, identifying the maximals of $\textbf{P}_{\boldsymbol{x}_{f},\phi}$ problems provides every information to construct the set $E^{*}(t_{f})$. This enables to utilize the existing results on time optimal problems to obtain useful conclusions on reachability sets. Sussmann (1995) uses PMP to find the maximals of minimum time problem (i.e., $\textbf{P}_{\boldsymbol{x}_{f},\phi}$ with $\phi<0$ described by Eqs. (8), (9), and (10)), and uses additional arguments (e.g., cycle argument) to find the structure of optimal solutions. Thus, the conclusions in Sussmann (1995) regarding maximals of minimum time problem will be leveraged throughout the following discussions to constuct $E^{*}(t_{f})$. However, we note that the other arguments that are not based on Eqs. (3) and (4), such as cycle argument, cannot be used for our reachability analysis.
@@ 3.2 Reachability of 3D curves with prescribed curvature bound
In this subsection, we present the reachability analysis of 3D curves with a prescribed curvature bound. The analysis is organized into two parts: first, we provide an implicit description of $\mathcal{G}_{3}(t_{f})$ by proving that every boundary point can be reached by curves belonging to the classes CCC, CSC, their subsegments, or H. Second, we present the complete description of $\pi_{3}\left(\mathcal{G}_{3}(t_{f})\right)$.
@@ 3.2.1 Implicit description of the reachability set
Proof. Proof of this theorem is largely an application of Theorem 3 to the system Eq. (9). We note that Eq. (9) satisfies Assumption 1 because it is a smooth control affine system with compact and convex restraint set, $\mathbb{B}^{3}$. (Recall that the maps $\boldsymbol{w}\mapsto\boldsymbol{y}\times\boldsymbol{w}$ correspond to skew-symmetric linear transformations of $\mathbb{R}^{3}$.)
[…]
@@
In terms of the notations used in the previous sections, $(\boldsymbol{\lambda},\boldsymbol{\mu})$ corresponds to the adjoint variable $-\boldsymbol{p}$ and $\nu$ corresponds to $-p_{0}\phi$. From here on, we will denote $\boldsymbol{\Xi}=(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\lambda},\boldsymbol{\mu},\nu,\boldsymbol{w})$ satisfying Eqs. (9) and (15) by lifted controlled trajectories, and lifted controlled arcs if it is defined on a compact interval. A lifted controlled trajectory is said to be $\mathcal{H}$-minimizing if $\mathcal{H}\left(\boldsymbol{\Xi}(t)\right)=\min\{\mathcal{H}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\lambda},\boldsymbol{\mu},\nu,\boldsymbol{v}):\boldsymbol{v}\in\mathbb{B}^{3}\}$ for a.e. $t$.
@@
    $\boldsymbol{\Xi}$ is $\mathcal{H}$-minimizing,
@@
These are merely a restatement of the PMP conditions of $\textbf{P}_{\boldsymbol{x}_{f},\phi}$ with arbitrary $\boldsymbol{x}_{f}$ and $\phi\neq 0$. Indeed, when studying the minimum time problems, or $\textbf{P}_{\boldsymbol{x}_{f},\phi}$ problems with $\phi<0$, in Sussmann (1995), $\nu\geq 0$ condition was present. This condition is omitted here because we consider $\phi>0$ case as well.
@@
Dependency of $\boldsymbol{\lambda}$ on $t$ is ignored since Eq. (15) implies that $\boldsymbol{\lambda}$ is constant in $t$. In Sussmann (1995), this formulation of the nontriviality condition—rather than the “naïve” version that requires $\left(\boldsymbol{\lambda},\boldsymbol{\mu}\right)\neq 0$—played a crucial role. This is because the maximum principle with the latter nontriviality condition provided no useful information, as every trajectory was maximal (Remark 3.2, Sussmann (1995)). Under the nontriviality condition in Eq. (16), the remaining part of the proof of this theorem demonstrates how maximum principle provides nontrivial conditions to sort out the maximals.
@@
Our proof follows the similar step, but with the three aforementioned modifications. For the rest of part 1, we introduce the major conclusions—primarily on the structure of nice trajectories—made in Sussmann (1995) that were solely grounded on the PMP conditions. As discussed in the previous subsection, these conclusions can be applied analogously to our proof. The summary is as follows:
@@
If $C\neq 0$, Eq. (22) can be modified in terms of $\tau$, which becomes exactly Eq. (11) with $\zeta=\frac{2\nu}{\sqrt{|C|}}$. Since nice trajectories are globally defined in this case, there is no need of further analysis how these connect up with trajectories such that $\boldsymbol{W}=0$. Additional remark on $C\neq 0$ case is that the condition of $\nu\geq 0$ was present in Sussmann (1995) because $\textbf{P}_{\boldsymbol{x}_{f},\phi}$ problems with $\phi<0$ were considered. Hence, the ‘H’ curves were confined to the ones with $\zeta\geq 0$ but here we allow $\zeta<0$ case as well.
@@
In this part, we study how nice trajectories connect up with trajectories such that $\boldsymbol{W}=0$ when it is not globally defined. This is essentially a process of identifying $S^{*}(t_{f})$ and $E^{*}(t_{f})$, but we alongside exclude some trajectories that have an alternative trajectory reaching the same endpoint through a controller that is not maximal.
Let us commence by considering a lifted controlled arc $\boldsymbol{\Xi}$ defined on an interval $I=[0,t_{f}]$ satisfying the PMP conditions. We define $L=\{t\in I:\boldsymbol{W}(t)\neq 0\}$ and its complement in $I$, $Q=I\setminus L=\{t\in I:\boldsymbol{W}(t)=0\}$. Since $\|\boldsymbol{W}\|=\|\boldsymbol{y}\times\boldsymbol{\mu}\|$ is continuous, its inverse image of the open set $\mathbb{R}\setminus\{0\}$, $L$, is open in $I$.
@@
First, if $\|\boldsymbol{\lambda}\|<|\nu|$, then Eq. (24) is a contradiction. Thus, $\boldsymbol{W}\neq 0$ and $\boldsymbol{\Xi}$ is a single piece of nice arc if it exists. If $\|\boldsymbol{\lambda}\|>|\nu|$, then $\dot{\boldsymbol{W}}(t_{1})\neq 0$ for all $t_{1}\in Q$. So each $t_{1}$ is an isolated point of $Q$ and $Q$ is a discrete set. This implies that $\boldsymbol{\Xi}$ consists of concatenation of number of nice trajectories that are connected up with each other at points such that $\boldsymbol{W}=0$. If $\|\boldsymbol{\lambda}\|=|\nu|$ then the zeros of $\boldsymbol{W}$ are not necessarily isolated and $\boldsymbol{\Xi}$ could consist of concatenations of nice trajectories with trajectories such that $\boldsymbol{W}\equiv 0$.
@@
Since $\boldsymbol{W}$ is differentiable, Taylor’s Theorem implies that for $t_{1}\in Q$, $\boldsymbol{W}(t)=\dot{\boldsymbol{W}}(t_{1})(t-t_{1})+o(|t-t_{1}|)$ and $\boldsymbol{w}(t)=sgn(t-t_{1})\frac{\dot{\boldsymbol{W}}(t_{1})}{\|\dot{\boldsymbol{W}}(t_{1})\|}+o(1)$ for $t\neq t_{1}$ near $t_{1}$. Therefore, the left and right side limits of $\boldsymbol{w}$ only change signs at $t_{1}$. This implies that the osculating plane of the two circular arcs being concatenated at $t_{1}$ coincides and their unit normal vectors are in the opposite directions. Hence, the state trajectory $\boldsymbol{x}(t)$ is a planar curve, concatenation of countable circular arcs. Moreover, $\boldsymbol{W}=0$ at each point of concatenation, so $\boldsymbol{\Xi}_{i}$ for $12\leq i\leq N-$ 1 correspond to maximal solutions of Eq. (22). Here, $\boldsymbol{\Xi}_{1}$ and $\boldsymbol{\Xi}_{N}$ are not necessarily maximal because $\boldsymbol{W}(0)$ and $\boldsymbol{W}(t_{f})$ are not necessarily 00.
@@
If $\nu=-\|\boldsymbol{\lambda}\|$, then it is easy to see that $\|\boldsymbol{W}\|=\|\boldsymbol{\lambda}\|\cos(t-t_{0})+\nu$ can be nonnegative only at isolated points. Hence, no lifted controlled trajectory $\boldsymbol{\Xi}$ on a nontrivial interval can exist if $\nu=-\|\boldsymbol{\lambda}\|$. If $\nu=\|\boldsymbol{\lambda}\|$, then maximal solutions of Eq. (22) are defined on open intervals of length $2\pi$. Since $L$ is open in $I$, it can be expressed as $L=\bigcup\limits_{J\in\mathcal{J}}J$ where $\mathcal{J}$ is a countable set of disjoint open subintervals of $I$. Then for any $J\in\mathcal{J}$ that does not contain 00 and $t_{f}$, $\boldsymbol{W}=0$ at its boundary points. Thus, such $J$ has a length of $2\pi$ and corresponds to the domain of a maximal nice trajectory. These maximal nice trajectories correspond to a single cycle. If the domain contains 00 or $t_{f}$, then corresponding $\varphi$ does not have to be maximal, and naturally, such nice trajectory is a circular arc of length $\leq 2\pi$.